How do I write the equation of a line that passes through the points (-1,4)and (-6,-2).
Answer by stanbon(75887)

write the equation of a line that passes through the points (-1,4)and (-6,-2).
-----
slope = (4--2)/(-1--6) = 6/5
-----------------
Form:: y = mx + b
Solve for "b"::
4 = (6/5)(-1) + b
b = 4+(6/5) = 26/5
-------
Equation:
y = (6/5)x + (26/5)
-------------------
Cheers,
Stan H.
-------